A collection of 15 coins is worth $1.00. If the coins are dimes and nickels, how many of each coin are there?
I have only been taught to do it when you are told that one of the coins has more than the other. Not sure how to set this up

Answer by MathTherapy(10552)


A collection of 15 coins is worth $1.00. If the coins are dimes and nickels, how many of each coin are there?
I have only been taught to do it when you are told that one of the coins has more than the other. Not sure how to set this up
Method 1: Solving for 2 unknowns
Let number of dimes and number of nickels, be D and N, respectively. Then we have:
D + N = 15 --------- Denomination-number equation
.1D + .05N = 1 --------- Denomination-value equation
Solve for D or N and substitute solved variable into either equation to determine value of other variable
Method 2: Solving for 1 unknown
Let number of dimes, be D
Then number of nickels = 15 - D. We now have:
.1D + .05(15 - D) = 1 -------- Value equation
Solve for D, and then SUBTRACT that value from 15 (total number of coins) to determine number of nickels
Then do a check!!
===================
If you need a complete and detailed solution, let me know!!
Send comments, “thank-yous,” and inquiries to “D” at MathMadEzy@aol.com.
Further help is available, online or in-person, for a fee, obviously.
For FREE info and answers to questions about the ASVAB exam, the NYS 3 – 8 city/state wide exams,GENERAL
MATH and HOMEWORK QUESTIONS, or MATH QUESTIONS related to the Regents Integrated Algebra,
Regents Geometry, Regents Algebra 2/Trigonometry, SHSAT, COOP/HSPT/TACHS, PSAT, SAT, ACT, SSAT/ISEE,
GRE, CLEP, and the TASC/GED, you can visit: http://asvabstudyzone.freeforums.net/.